# LRPO: Enhancing Blind Face Restoration through Online Reinforcement Learning

### v1

@@ 5.1 Experimental Settings
Training and Testing Data. We only use 3000 face images from the FFHQ (Karras et al., 2021) dataset for training. The degradation strategy from HQ to LQ is based on the following degradation function: $\mathbf{I}_{\text{LQ}}=\left\{\left[\left(\mathbf{I}_{\text{HQ}}\otimes\boldsymbol{k}_{\sigma}\right)_{\downarrow_{r}}+\boldsymbol{n}_{\delta}\right]_{\mathrm{JPEG}_{q}}\right\}_{\uparrow_{r}},$ where the HQ images are first convolved with a Gaussian kernel $\boldsymbol{k}_{\sigma}$, followed by a downsampling with a factor of $r$, and then corrupted with Gaussian noise $\boldsymbol{n}_{\delta}$. Subsequently, the images undergo JPEG compression with a quality factor of $q$. Finally, the LQ image is resized back to the original $512\times 512$. Here, $\sigma$, $r$, $\delta$, and $q$ are randomly sampled from the intervals $[0.1,12]$, $[1,12]$, $[0,15]$, and $[30,100]$, respectively. Follow previous work Wang et al. (2021); Gu et al. (2022), we employ the synthetic dataset CelebA-Test (Karras et al., 2017) and two real-world datasets (Wang et al., 2021) (i.e., LFW-Test and WebPhoto-Test) to validate our method.
Evaluation Metrics. On the Celeba-Test dataset, we utilized six common reference-based metrics: SSIM (Wang et al., 2004), PSNR, LPIPS (Zhang et al., 2018), CLIP Score (Hessel et al., 2021), Deg. (Wang et al., 2021), and LMD (Gu et al., 2022), where Deg. and LMD are identity consistency metrics, along with four non-reference metrics: MUSIQ (Ke et al., 2021), MANIQA (Yang et al., 2022), CLIPIQA (Wang et al., 2023a), and Aesthetic (LAION-AI, 2022).
Comparison Methods. We compare with not only the base models but also the latest state-of-the-art methods, including GFPGAN (Chan et al., 2021), CodeFormer (Zhou et al., 2022), VQFR (Gu et al., 2022), DR2+SPAR (Wang et al., 2023b), RestoreFormer (Wang et al., 2022), DifFace (Yue & Loy, 2024), OSEDiff (Wu et al., 2024), DiffBIR (Lin et al., 2024), FlowIE (Zhu et al., 2024) and InterLCM (Li et al., 2025).
@@
5.2 Main Results
Results on Synthetic Data. As shown in Table 1, LRPO achieves improvements on all metrics compared with DiffBIR on the synthetic CelebA-Test dataset. These results indicate that our RL framework simultaneously improves perceptual quality and identity preservation in restored faces. Furthermore, LRPO achieves superior performance compared to state-of-the-art approaches across the majority of evaluation metrics, including SSIM, LMD, and MUSIQ, confirming that it enhances identity consistency while maintaining perceptual quality. Figure 3 demonstrates LRPO’s superior performance over methods that fail to restore faces satisfactorily. LRPO delivers more realistic textures than the baseline, better identity alignment than DR2 and OSEDiff, and more natural results without the over-smoothing seen in InterICM and FlowIE.
@@ 5.2 Main Results
5.3 Ablation Study
We conduct the ablation study on CelebA-Test dataset. As shown in Table 4, we analyze the effects of four key components: Reinforcement Learning (RL), Kullback-Leibler divergence (KL), GT guided likelihood regularization (Reg.), and noise-level advantage assignment (AdA). Variant 1 demonstrates improvements across all metrics after incorporating RL, confirming that RL directly enhances BFR performance. However, adding KL divergence in Variant 2 degrades all metrics without improving visual quality (See Figure 5(a)). We therefore remove KL divergence from the RL objective in Eq. LABEL:eq:lrpo, reducing computational cost while maintaining visual quality. Without AdA (Variant 3), the model suffers from detail blurring caused by over-optimization during late denoising stages (Figure 5(b)). Removing Reg. (Variant 4) leads to decreased SSIM scores and poor fidelity, with the model producing unrealistic, fantasy-like textures as shown in Figure 5(c). Training dynamics show that our noise-level advantage assignment facilitates faster convergence to high-reward restoration trajectories, while GT guided likelihood regularization enhances structural fidelity (Appendix D).
@@ 5.3 Ablation Study
Figure 5: Ablation study visualizations.
@@
Limitation. While our method surpasses existing approaches, certain failure cases remain. Figure 6 demonstrates that rare specialized and individualized objects (e.g., jewelry) are restored with artifacts due to limited prior knowledge in the base model’s training data. These limitations may require stronger foundation models or more comprehensive training datasets.
@@ Appendix C Composite Reward Function Details
Human Preference Reward ($r_{\text{pref}}$). The human preference evaluation we use is based on the previous work (Wu et al., 2025), called the Face Reward Model. Trained on several human preference datasets, it is able to provide face restoration evaluations with high human consistency. The Face Reward Model’s input requires both $\hat{\bm{x}}_{0}$ and $\bm{c}_{\text{text}}$, the latter corresponding to the GT face.
Perceptual Quality Reward ($r_{\text{aq}}$). This component, $r_{\text{aq}}$, offers a no-reference evaluation of the image’s absolute quality. We employ the CLIPIQA (CLIP-based Image Quality Assessment) metric Wang et al. (2023a) from the pyiqa library. It assesses overall perceptual quality and realism without requiring a reference image, making it effective for identifying artifacts.
@@
LPIPS. The Learned Perceptual Image Patch Similarity (LPIPS) (Zhang et al., 2018) metric computes the distance between two images in a deep feature space, which correlates well with perception. As LPIPS is a distance metric (lower is better), we convert it into a similarity reward (higher is better) via the transformation: $r_{\text{lpips}}=1.0-\text{LPIPS}(\hat{\bm{x}}_{0},\bm{x}_{\text{GT}})$.

### v2

@@ 5.1 Experimental Settings
Training and Testing Data. We use face images from the FFHQ [16] dataset for training. The degradation strategy from HQ to LQ is based on the following degradation function: $\mathbf{I}_{\text{LQ}}=\left\{\left[\left(\mathbf{I}_{\text{HQ}}\otimes\boldsymbol{k}_{\sigma}\right)_{\downarrow_{r}}+\boldsymbol{n}_{\delta}\right]_{\mathrm{JPEG}_{q}}\right\}_{\uparrow_{r}},$ where the HQ images are first convolved with a Gaussian kernel $\boldsymbol{k}_{\sigma}$, followed by a downsampling with a factor of $r$, and then corrupted with Gaussian noise $\boldsymbol{n}_{\delta}$. Subsequently, the images undergo JPEG compression with a quality factor of $q$. Finally, the LQ image is resized back to the original $512\times 512$. Here, $\sigma$, $r$, $\delta$, and $q$ are randomly sampled from the intervals $[0.1,12]$, $[1,12]$, $[0,15]$, and $[30,100]$, respectively. Follow previous work Wang et al. [36], Gu et al. [9], we employ the synthetic dataset CelebA-Test [13] and two real-world datasets [36] (i.e., LFW-Test and WebPhoto-Test) to validate our method.
Evaluation Metrics. On the Celeba-Test dataset, we utilized six common reference-based metrics: SSIM [37], PSNR, LPIPS [50], CLIP Score [11], Deg. [36], and LMD [9], where Deg. and LMD are identity consistency metrics, along with four non-reference metrics: MUSIQ [17], MANIQA [43], CLIPIQA [34], and Aesthetic [19].
Comparison Methods. We compare with not only the base models but also the latest state-of-the-art methods, including GFPGAN [4], CodeFormer [52], VQFR [9], DR2+SPAR [39], RestoreFormer [38], DifFace [48], OSEDiff [41], DiffBIR [21], FlowIE [53] and InterLCM [20]. Among them, FlowIE and InterLCM use ODE solvers.
@@
5.2 Main Results
Results on Synthetic Data. As shown in Table 1, LRPO achieves improvements on all metrics compared with DiffBIR on the synthetic CelebA-Test dataset. These results indicate that our RL framework simultaneously improves perceptual quality and identity preservation in restored faces. Furthermore, LRPO achieves superior performance compared to state-of-the-art approaches across the majority of evaluation metrics, including SSIM, LMD, and MUSIQ, confirming that it enhances identity consistency while maintaining perceptual quality. Figure 3 demonstrates LRPO’s superior performance over methods that fail to restore faces. LRPO delivers more realistic textures than the baseline, better identity alignment than OSEDiff, and more natural results without the over-smoothing seen in OSEDiff and DiffBIR. Appendix F contains the complete qualitative results.
@@ 5.2 Main Results
5.3 Ablation Study
We conduct the ablation study on CelebA-Test dataset. As shown in Table 4, we analyze the effects of four key components: Reinforcement Learning (RL), Kullback-Leibler divergence (KL), GT guided likelihood regularization (Reg.), and noise-level advantage assignment (AdA). Variant 1 demonstrates improvements across all metrics after incorporating RL, confirming that RL directly enhances BFR performance. However, adding KL divergence in Variant 2 degrades all metrics without improving visual quality (See Figure 5(a)). We therefore remove KL divergence from the RL objective in Eq. 13, reducing computational cost while maintaining visual quality. Without AdA (Variant 3), the model suffers from detail blurring caused by over-optimization during late denoising stages (Figure 5(b)). Removing Reg. (Variant 4) leads to decreased SSIM scores and poor fidelity, with the model producing unrealistic, fantasy-like textures as shown in Figure 5(c). Training dynamics show that our noise-level advantage assignment facilitates faster convergence to high-reward restoration trajectories, while GT guided likelihood regularization enhances the consistency with the true distribution. (Appendix D).
Figure 5: Ablation study visualizations.
@@ 5.3 Ablation Study
In addition, Figure 6 shows the experiments conducted to select the number of trajectories $G$ within each group. Because $G=9$ exhibits superior performance in terms of both structural similarity and identity consistency, we set $G=9$ for the training stage. The proposed likelihood regularizer is applied only during the last $S$ denoising steps. Throughout training, varying $S$ leads to different reward curves, as shown in Figure 7. Choosing an excessively large $S$ introduces overly strong penalties that conflict with the RL learning trajectory, ultimately causing training collapse. $S=5$ achieves convergence to a higher reward score without causing training collapse; therefore, we adopt $S=5$.
Figure 6: Quantitative metrics under different group sizes $G$.
Figure 7: Reward curves of different $S$ values within the likelihood regularization term
@@
Supplementary Material
Enhancing Blind Face Restoration through Online Reinforcement Learning
@@ Appendix C Composite Reward Function Details
Human Preference Reward ($r_{\text{pref}}$). The human preference evaluation we use is based on the previous work [40], called the Face Reward Model. Trained on several human preference datasets, it is able to provide face restoration evaluations with high human consistency. The Face Reward Model’s input requires both $\hat{\bm{x}}_{0}$ and $\bm{c}_{\text{text}}$, the latter corresponding to the GT face.
Perceptual Quality Reward ($r_{\text{aq}}$). This component, $r_{\text{aq}}$, offers a no-reference evaluation of the image’s absolute quality. We employ the CLIPIQA (CLIP-based Image Quality Assessment) metric [34] from the pyiqa library. It assesses overall perceptual quality and realism without requiring a reference image, making it effective for identifying artifacts.
@@
LPIPS. The Learned Perceptual Image Patch Similarity (LPIPS) [50] metric computes the distance between two images in a deep feature space, which correlates well with perception. As LPIPS is a distance metric (lower is better), we convert it into a similarity reward (higher is better) via the transformation: $r_{\text{lpips}}=1.0-\text{LPIPS}(\hat{\bm{x}}_{0},\bm{x}_{\text{GT}})$.
@@ Appendix F More Qualitative Results
Appendix G Limitation.
Despite the improvements over existing approaches, our method is not without limitations. One failure case involves the restoration of highly individualized attributes, such as complex jewelry and accessories (Figure 10). Due to the limited representation of such rare objects in the training data of the base diffusion model, our method may fail to hallucinate plausible high-frequency details for these regions, occasionally resulting in blurred or distorted artifacts. Additionally, while the inference speed remains consistent with the base model, the training stage of LRPO requires additional computational resources for online sample generation and reward calculation compared to standard supervised fine-tuning. Future work aims to mitigate the artifact issue by exploring more diverse datasets and improving the efficiency of the RL alignment process.